a + 9b=4
a + 4b= -1

Found 2 solutions by checkley79, Alan3354:
Answer by checkley79(3341)

a + 9b=4
a + 4b= -1 SUBTRACT.
-----------------------------
5b=5
b=5/5
b=1 ans.
a+9*1=4
a+9=4
a=4-9
a=-5 ans.
proof"
-5+4*1=-1
-5+4=-1
-1=-1

Answer by Alan3354(69443)